A body weighing 40 lbs. starts from rest and slides down a plane at an angle of 30ᵒ with the horizontal for which the coefficient of static friction is 0.30. How far will it move during the third second?


1
Expert's answer
2021-05-20T10:10:26-0400
"ma=mg(\sin{30}-\mu \cos{30})\\a=9.8(\sin{30}-0.3 \cos{30})=2.35\frac{m}{s^2}"

"d=0.5(2.35)(3^2-2^2)=5.9\ m"